OpenStudy (anonymous):

Determine whether the formula describes y as a function of x. y = 9x + 1

OpenStudy (anonymous):

Easy x - 3y = 9 x = 9 + 3y 3y = x - 9 y = (x-9)/3 --> TA DAH !! y is a function of x !

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Welcome(: